OpenStudy (animelover91804):

Express the rate as a unit rate: $258 for 6 tires.

OpenStudy (anonymous):

Wouldn't you divide?

OpenStudy (animelover91804):

Idk... Maybe?

OpenStudy (anonymous):

256/6=42.66666666666667 change that into a fraction :D

OpenStudy (animelover91804):

Oh....

OpenStudy (anonymous):

6...... maybye? ;D

OpenStudy (johnweldon1993):

Correct, the unit rate means how much would it be for just 1 So if its 258 for 6....how much for just 1?

OpenStudy (anonymous):

43 i think

OpenStudy (anonymous):

@johnweldon1993 is that right?

OpenStudy (johnweldon1993):

Correct....so the unit rate is \(\large $43\) per tire

OpenStudy (anonymous):

WOOP totally just did that ;D

OpenStudy (anonymous):

congrats XD

OpenStudy (anonymous):

Ty ty *applause*

OpenStudy (anonymous):

*throws roses* -cheering-

OpenStudy (animelover91804):

XD Thanks

OpenStudy (anonymous):

Ikr I'm magnificent right :D

OpenStudy (animelover91804):

Can you help me with a few more?

OpenStudy (anonymous):

suuuuuure xD

OpenStudy (anonymous):

just tag me in them and my geniusness will...uhhh yea ;D

OpenStudy (anonymous):

XDDDD

OpenStudy (animelover91804):

Solve the proportion using cross product property\[\frac{ 72 }{ 120 } = \frac{ 9 }{ y }\]

OpenStudy (anonymous):

y=15

OpenStudy (animelover91804):

A 2006 Harley-Davidson Fat-Boy Motorcycle can travel 69 miles on 1 gallons of gasoline. Write and solve a proportion to determine how many gallons of gasoline this motorcycle will need to travel 350 miles. (Round answer to nearest tenth.)

OpenStudy (anonymous):

5.07246 so round that it'd be 5

OpenStudy (animelover91804):

A 32 ounce bottle of fresh-squeezed fruit juice sells for $3.29. A 24-ounce bottle of the same juice sells for $1.97. Which bottle is the better deal? * Show your work to PROVE which is the better deal by finding the cost per ounce.

OpenStudy (anonymous):



OpenStudy (anonymous):

sorry that took so long and your choice would be the 24 ounce

OpenStudy (animelover91804):

Mr. Lenz rented a scooter during her beach vacation. The cost of the scooter rental is shown for different periods of time. What does the point (1, 10) represent? Choose the correct answer below. * a. The unit rate of $10 per hour. b. The unit rate of $1 per hour. c. The cost of renting the scooter for 10 hours is $1. d. The cost of renting the scooter for 10 hours is $10.

OpenStudy (anonymous):

a